okay Mrs. Brook lacks 7 years from being five times as old as her son. Six years from now she will lack 3 years from being three times as old as her son is then. Find each of the numbers
Answer by stanbon(75887)

Mrs. Brook lacks 7 years from being five times as old as her son. Six years from now she will lack 3 years from being three times as old as her son is then. Find each of the numbers.
-----------------------
Equations:
b = 5s-7
(b+6) = 3(s+6)-3
-----
Substitute for "b" and solve for "s":
5s-7+6 = 3s+18-3
2s = 16
s = 8 yrs old (her son's age now)
---
Since b = 5s-7, b = 5*8-7 = 33 yrs old (Mrs. Brook's age now)
===================
Cheers,
Stan H.
---------